### v1

@@ Appendix C More Experiment Results
In this section, we show more experiment results. Section C.1 shows more experiment results on much larger graphs, where the AI-inspired methods cannot handle. Section C.2 show the serialization results on more graphs. Section C.3 show a more detailed results between LwD and Deg-Greedy, which applies degree-based serialization as a subprocedure. Section C.4 shows the full results when adding local search as a post-processing procedure. Finally, Section C.5 discusses more results on the ratio of the solution found by different algorithms to $\frac{n\ln d}{d}$ on ER graphs, which relates to the theoretical conjecture mentioned in Section 4.4.
C.1 Larger graphs
@@
C.2 Serialization
Figures 4 and 5 shows the percentage to choose the smallest possible degree node on different part of the serialization across various algorithms for all ER graphs and all BA graphs with nodes $n\leq 3000$, respectively. The serialization process is discussed in Section 4.2. Missing bars are algorithms which we do not get results due to computational limit, same as in Tables 1 and 2.
@@ C.2 Serialization
We also note that BA graphs $(n=300,m=50)$, $(n=1000,m=150)$, and $(n=3000,m=500)$ are outliers. Most algorithms have percentage close to $100\%$ for all parts. This is because these graphs are rather different from other BA graphs. They have an easily found large MIS, which is the $m$ nodes initially in the graph at the start of the BA generation process. (Albert & Barabási, 2002). From Table 2, we can see many algorithms can find these MIS and report a MIS size of $m$ for these graphs. This suggests that for this special type of BA graphs, our serialization analysis can observe different characteristics from other BA graphs.
C.3 Comparison between Deg-Greedy and LwD
@@
The heatmaps for overall percentages suggest that LwD is not very similar to Deg-Greedy like LTFT. From the heatmaps for different 1/3 parts, we can see the percentage increases from the 1st 1/3 to the 3rd 1/3 for all the datasets (different parameters of the synthetic graphs). This aligns with our “counterfactual” serialization results for LwD in Section 4.2, where we also observe the percentage increases clearly from 1st 1/3 to 3rd 1/3. This shows that our serialization method in Section 4.2 can reflect the pattern correctly.
The percentages here in the heatmap is smaller than the percentages for “counterfactual” serialization in the bar graphs in Section 4.2. This is likely due to the fact that in the “counterfactual” serialization we repeat the serialization process for $100$ times and report the highest percentages we get, while here we only do serialization once for each step.
C.4 Local search
@@
Heuristics GPU-acc Learning-based param Deg-Greedy OnlineMIS ReduMIS iSCO PCQO LwD LTFT DIFUSCO ER Graphs 100,10 29.50 (0.25) 30.50 30.50 30.62 (0) 27.00 (2) 30.38 (0) 29.00 (0.38) 30.25 (0) 100,30 13.62 (0) 14.00 14.75 14.50 (0) 12.00 (0.50) 14.38 (0) 13.25 (0.13) 13.88 (0) 300,10 92.12 (0.50) 93.88 94.38 94.75 (0) 91.50 (0.88) 94.25 (0) 90.12 (1.50) 93.50 (0) 300,30 44.75 (0.25) 47.88 47.88 47.62 (0) 43.00 (1.25) 46.88 (0) 44.00 (0.75) 44.62 (0.74) 300,100 16.12 (0) 18.00 18.38 18.00 (0) 14.25 (0.87) 17.12 (0.12) 16.25 (0) 16.88 (0.26) 1000,10 305.38 (2.13) 314.75 316.13 315.62 (0) 308.12 (3) 311.25 (0) 300.00 (3.00) 306.38 (2.50) 1000,30 152.75 (1.75) 158.88 163.75 163.50 (0) 148.50 (5.50) 158.62 (0.24) 151.38 (1.38) 150.62 (6.87) 1000,100 60.75 (0.13) 64.75 66.63 66.50 (0) 59.00 (3.38) 64.12 (0.24) 61.88 (1.00) 57.88 (2.50) 1000,300 22.25 (0) 25.00 25.75 24.62 (0) 21.25 (1.25) 20.75 (1.63) 22.62 (0) 21.62 (0.74) 3000,10 913.62 (6.50) 947.25 954.25 950.88 (0) 927.00 (8.00) 935.38 (1.26) 900.62 (12.37) 918.00 (16) 3000,30 456.12 (4.24) 480.88 493.13 491.62 (0) 453.88 (11.50) 474.00 (0.75) 454.38 (5.38) 442.75 (29.37) 3000,100 185.62 (2) 194.38 201.50 200.38 (0) 180.25 (9.87) 191.50 (0.75) 185.62 (1.62) 182.25 (10.87) 3000,300 74.00 (0.50) 77.63 80.75 78.88 (0) 68.00 (4.50) 74.38 (0.50) 3000,1000 23.38 (0) 26.00 26.25 21.38 (1.50) 23.62 (0) BA Graphs 100,5 39.25 (0) 39.50 39.50 39.50 (0) 36.62 (2.74) 39.50 (0) 38.50 (0.38) 39.38 (0) 100,15 21.00 (0) 21.63 21.63 21.62 (0) 18.88 (0.50) 21.62 (0) 20.62 (0) 21.25 (0) 300,5 122.38 (0.26) 123.13 123.13 123.12 (0) 120.25 (2.63) 123.12 (0) 118.62 (3.00) 123.00 (0.12) 300,15 70.00 (0.75) 71.38 71.38 71.38 (0) 67.62 (2.00) 70.75 (0) 66.62 (1.87) 70.00 (0.50) 300,50 39.25 (1.75) 49.88 50.00 50.00 (0) 32.38 (5.13) 50.00 (0) 43.62 (1.87) 50.00 (0) 1000,5 413.38 (2) 417.13 417.13 417.12 (0) 412.62 (1.74) 416.00 (0) 400.25 (14.50) 417.12 (0) 1000,15 234.88 (1.63) 245.00 246.38 246.25 (0) 235.50 (4.50) 243.12 (0) 230.50 (6.25) 237.88 (1.50) 1000,50 108.38 (1.00) 115.75 116.88 116.75 (0) 110.25 (2.37) 113.12 (0.12) 106.75 (0.37) 108.75 (3.50) 1000,150 90.88 (7.63) 150.00 150.00 150.00 (0) 112.75 (3.13) 150.00 (0) 87.62 (5.37) 3000,5 1241.50 (4.75) 1257.00 1257.13 1255.62 (0) 1242.12 (3.37) 1248.50 (0.38) 1213.12 (35.87) 1254.75 (0.37) 3000,15 714.62 (7.50) 749.63 754.50 752.00 (0) 723.50 (12.88) 730.50 (2.75) 693.00 (31.12) 731.75 (5.37) 3000,50 339.00 (3.62) 362.63 369.75 368.25 (0) 348.50 (9.62) 341.00 (4.12) 334.00 (10.12) 3000,150 142.88 (2.38) 160.25 165.75 164.00 (0) 150.50 (6.50) 146.50 (1.75) 3000,500 172.12 (9.12) 500.00 500.00 500.00 (0) 229.62 (5.74)
@@
Table 8 shows the full results after incorporating $2$-improvement local search from the ARW local search algorithm (Andrade et al., 2012) as a post-processing step, which is discussed in Section 4.3.
C.5 More results on the ratio
In addition to what we show in Section 4.4, Figure 7 shows the ratio of MIS size to $\frac{n\ln d}{d}$ for ER graphs with number of nodes $n$ and average degree $d$ across more algorithms. We can see that ReduMIS, OnlineMIS, and iSCO has consitently high ratios more than $1.2$. Ran-Greedy stays around $1.0$ for all $(n,d)$. Other algorithms, including Deg-Greedy, all have higher ratios for sparser graphs, but lower ratios (close to $1$) for denser graphs.

### v2

@@ Appendix C More Experiment Results
In this section, we show more experiment results. Section C.1 shows the performance of various algorithms as in Section 4 on other types graphs.Section C.2 shows more experiment results on much larger graphs, where the AI-inspired methods cannot handle. Section C.4 show the serialization results on more graphs. Section C.5 show a more detailed results between LwD and Deg-Greedy, which applies degree-based serialization as a subprocedure. Section C.6 shows the full results when adding local search as a post-processing procedure.
C.1 Results for BA graphs, RB graphs, and real-world graphs
In this section we report the performance of various algorithms on Barabási–Albert (BA) graphs in Table 5, RB graphs in Table 6, and real-world graphs in Table 7.
[…]
Heuristics GPU-acc Learning-based $n$ $m$ Deg-Greedy OnlineMIS ReduMIS iSCO PCQO LwD LTFT DIFUSCO DiffUCO 100 5 39.50 39.50 39.50 39.50 39.50 39.50 38.12 39.38 39.25 15 21.00 21.63 21.63 21.62 21.63 21.62 20.62 21.25 21.25 300 5 121.88 123.13 123.13 123.12 123.00 123.12 115.62 122.88 123.00 15 69.38 71.38 71.38 71.38 71.25 70.75 64.75 69.50 69.25 50 38.00 49.88 50.00 50.00 50.00 50.00 41.75 50.00 50.00 1000 5 412.63 417.13 417.13 417.12 415.75 416.00 385.75 417.12 416.00 15 232.75 245.00 246.38 246.25 242.00 243.12 224.25 236.38 240.75 50 107.50 115.75 116.88 116.75 113.50 113.00* 106.38 105.25 56.38† 150 81.50 150.00 150.00 150.00 150.00 150.00* 82.25 135.00† 3000 5 1236.63 1257.00 1257.13 1255.62 1243.00 1248.12 1177.25 1254.38 1252.63† 15 709.50 749.63 754.50 752.00 722.25 727.75* 661.88 726.38 737.38† 50 335.38 362.63 369.75 368.25 357.13 336.88* 323.88 142.25† 150 147.25 160.25 165.75 164.00 152.75 144.75 500 141.00 500.00 500.00 491.00 223.88 10000 5 4128.75 4206.00 4205.38 4190.47 3640.75 15 2377.88 2525.00 2534.00 2517.07 1946.38 50 1122.13 1228.88 1251.13 1241.98 938.38 150 511.25 555.75 580.13 573.37 433.88 500 192.50 209.13 217.50 150.63 1500 67.25 1500.00 1500.00
Table 6: (Comparison of the performance of different algorithms on RB graphs) Table 1 on RB graphs [47]. Datasets are from [21]. The numbers within ±$\pm 1\%$ of the best are highlighted.
Dataset no. of nodes Heuristics GPU-acc Learning-based Deg-Greedy OnlineMIS ReduMIS iSCO PCQO LwD LTFT DIFUSCO DiffUCO RB-small 200-300 19.11 20.11 20.14 20.05 20.08 16.44 18.63 18.23 19.26 RB-large 800-1200 39.16 42.66 42.95 41.53 39.62 32.55 38.32 36.22 38.91
C.2 Larger graphs
@@
C.3 Comparison between Deg-Greedy and LTFT
We conducted the exact same analysis as in Section 5.1 for BA graphs. The results are in Figure 2.
Figure 2: Percentage of rounds when LTFT selects the node with smallest possible degree on BA graphs. On BA graphs, the percentage to choose the smallest possible node is generally lower, but on denser BA graphs LTFT also behaves more like degree-based greedy.
C.4 Serialization
Figures 3 and 4 shows the percentage to choose the smallest possible degree node on different part of the serialization across various algorithms for all ER graphs and all BA graphs with nodes $n\leq 3000$, respectively. The serialization process is discussed in Section 5.2. Missing bars are algorithms which we do not get results due to computational limit, same as in Tables 1 and 5.
@@ C.4 Serialization
We also note that BA graphs $(n=300,m=50)$, $(n=1000,m=150)$, and $(n=3000,m=500)$ are outliers. Most algorithms have percentage close to $100\%$ for all parts. This is because these graphs are rather different from other BA graphs. They have an easily found large MIS, which is the $m$ nodes initially in the graph at the start of the BA generation process. [46]. From Table 5, we can see many algorithms can find these MIS and report a MIS size of $m$ for these graphs. This suggests that for this special type of BA graphs, our serialization analysis can observe different characteristics from other BA graphs.
C.5 Comparison between Deg-Greedy and LwD
@@
The heatmaps for overall percentages suggest that LwD is not very similar to Deg-Greedy like LTFT. From the heatmaps for different 1/3 parts, we can see the percentage increases from the 1st 1/3 to the 3rd 1/3 for all the datasets (different parameters of the synthetic graphs). This aligns with our “counterfactual" serialization results for LwD in Section 5.2, where we also observe the percentage increases clearly from 1st 1/3 to 3rd 1/3. This shows that our serialization method in Section 5.2 can reflect the pattern correctly.
The percentages here in the heatmap is smaller than the percentages for “counterfactual" serialization in the bar graphs in Section 5.2. This is likely due to the fact that in the “counterfactual" serialization we repeat the serialization process for $100$ times and report the highest percentages we get, while here we only do serialization once for each step.
C.6 Local search
@@
Heuristics GPU-acc Learning-based param Deg-Greedy OnlineMIS ReduMIS iSCO PCQO LwD LTFT DIFUSCO ER Graphs 100,10 29.50 (0.25) 30.50 30.50 30.62 (0) 30.63 (0) 30.38 (0) 29.00 (0.38) 30.25 (0) 100,30 13.62 (0) 14.00 14.75 14.50 (0) 14.50 (0) 14.38 (0) 13.25 (0.13) 13.88 (0) 300,10 92.12 (0.50) 93.88 94.38 94.75 (0) 94.63 (0) 94.25 (0) 90.12 (1.50) 93.50 (0) 300,30 44.75 (0.25) 47.88 47.88 47.62 (0) 47.63 (0) 46.88 (0) 44.00 (0.75) 44.62 (0.74) 300,100 16.12 (0) 18.00 18.38 18.00 (0) 18.00 (0) 17.12 (0.12) 16.25 (0) 16.88 (0.26) 1000,10 305.38 (2.13) 314.75 316.13 315.62 (0) 310.75 (0.75) 311.25 (0) 300.00 (3.00) 306.38 (2.50) 1000,30 152.75 (1.75) 158.88 163.75 163.50 (0) 159.13 (0.50) 158.62 (0.24) 151.38 (1.38) 150.62 (6.87) 1000,100 60.75 (0.13) 64.75 66.63 66.50 (0) 60.88 (0.75) 64.12 (0.24) 61.88 (1.00) 57.88 (2.50) 1000,300 22.25 (0) 25.00 25.75 24.62 (0) 23.25 (0) 20.75 (1.63) 22.62 (0) 21.62 (0.74) 3000,10 913.62 (6.50) 947.25 954.25 950.88 (0) 927.38 (4.13) 935.38 (1.26) 900.62 (12.37) 918.00 (16) 3000,30 456.12 (4.24) 480.88 493.13 491.62 (0) 469.75 (5.50) 474.00 (0.75) 454.38 (5.38) 442.75 (29.37) 3000,100 185.62 (2) 194.38 201.50 200.38 (0) 189.75 (4.12) 191.50 (0.75) 185.62 (1.62) 182.25 (10.87) 3000,300 74.00 (0.50) 77.63 80.75 78.88 (0) 70.88 (1.63) 74.38 (0.50) 3000,1000 23.38 (0) 26.00 26.25 23.00 (0) 23.62 (0) BA Graphs 100,5 39.25 (0) 39.50 39.50 39.50 (0) 39.50 (0) 39.50 (0) 38.50 (0.38) 39.38 (0) 100,15 21.00 (0) 21.63 21.63 21.62 (0) 21.63 (0) 21.62 (0) 20.62 (0) 21.25 (0) 300,5 122.38 (0.26) 123.13 123.13 123.12 (0) 123.00 (0) 123.12 (0) 118.62 (3.00) 123.00 (0.12) 300,15 70.00 (0.75) 71.38 71.38 71.38 (0) 71.25 (0) 70.75 (0) 66.62 (1.87) 70.00 (0.50) 300,50 39.25 (1.75) 49.88 50.00 50.00 (0) 50.00 (0) 50.00 (0) 43.62 (1.87) 50.00 (0) 1000,5 413.38 (2) 417.13 417.13 417.12 (0) 415.88 (0.13) 416.00 (0) 400.25 (14.50) 417.12 (0) 1000,15 234.88 (1.63) 245.00 246.38 246.25 (0) 242.12 (0.12) 243.12 (0) 230.50 (6.25) 237.88 (1.50) 1000,50 108.38 (1.00) 115.75 116.88 116.75 (0) 114.00 (0.50) 113.12 (0.12) 106.75 (0.37) 108.75 (3.50) 1000,150 90.88 (7.63) 150.00 150.00 150.00 (0) 150.00 (0) 150.00 (0) 87.62 (5.37) 3000,5 1241.50 (4.75) 1257.00 1257.13 1255.62 (0) 1245.38 (2.38) 1248.50 (0.38) 1213.12 (35.87) 1254.75 (0.37) 3000,15 714.62 (7.50) 749.63 754.50 752.00 (0) 729.38 (7.13) 730.50 (2.75) 693.00 (31.12) 731.75 (5.37) 3000,50 339.00 (3.62) 362.63 369.75 368.25 (0) 359.25 (2.13) 341.00 (4.12) 334.00 (10.12) 3000,150 142.88 (2.38) 160.25 165.75 164.00 (0) 155.38 (2.63) 146.50 (1.75) 3000,500 172.12 (9.12) 500.00 500.00 491.00 (0) 229.62 (5.74)
@@
Table 9 shows the full results after incorporating $2$-improvement local search from the ARW local search algorithm [33] as a post-processing step, which is discussed in Section 5.3.
Appendix D Perspective for theoreticians: Does KaMIS refutes famous conjecture?
For theoretical analysis for MIS on ER graphs and regular graphs, see [51, 84, 54, 49]. In ER graphs with $n$ nodes and average degree $d$, the MIS has size $\frac{2n\ln d}{d}$ for asymptotically large $n$ and $d$, and simplest random greedy achieves half-optimal at $\frac{n\ln d}{d}$ [52]. Yet, there is no known polynomial-time algorithm which can achieve MIS size $(1+{\epsilon})\frac{n\ln d}{d}$ for any constant ${\epsilon}$ and it is conjectured that polynomial-time algorithms cannot do better than $(1+o(1))\frac{n\ln d}{d}$ [51]. Thus it is natural to measure the goodness of an algorithm by the ratio of the MIS size obtained to $\frac{n\ln d}{d}$. This ratio can faciliate comparison across different $n$ and $d$’s.
Figure 6 shows the ratio of MIS size to $\frac{n\ln d}{d}$ for ER graphs with number of nodes $n$ and average degree $d$ across various algorithms. Surprisingly for ReduMIS and OnlineMIS this ratio is consistently larger than $1.2$ and often larger than $1.3$ even for fairly large $n,d$. These results seem to bring the conjecture in doubt, but not disprove per se since the conjecture concerns asymptotically large $n$ and $d$. But our results could encourage further analysis and potential collaboration between researchers on theoretical and empirical aspects of MIS on random graphs.